# IMU: Influence-guided Machine Unlearning

### v2

@@ 7.3 Experiments on SalUn
[…]
Table 9: Unlearning performance of SalUn under the same settings.
7.4 Changes of influence score over epochs
In Figure 5, we respectively present the changes in the influence scores of $100$ samples randomly selected from ${\mathcal{D}}_{f}$ over the $5$ unlearning epochs in the class-wise and sample-wise forgetting tasks. In the class-wise forgetting task, since the samples are all belong to the same class, their values should all be negative according to the definition of the influence function formula. It is verified in Figure 5, and we can see the changes in the influence score of a specific sample after epoch $2$ do not significantly fluctuate. Additionally, compared to the class-wise task, since the samples are randomly selected from different classes in the sample-wise forgetting scenario, the overall influence of each sample on forgetting data is possibly positive or negative. We can also observe that with more forgetting rounds, the number of samples with negative influence values gradually increases, demonstrating the effectiveness of our unlearning method.
Figure 5: Change of sample influences for 55 epochs in CIFAR-10, the left column represents the class-wise forgetting tasks and the right column represents the random-wise forgetting tasks.
@@ 7.4 Changes of influence score over epochs
Table 10: Examples of responses after unlearning on TOFU (Forget05) against QAs targeted for unlearning. The bold font highlights the key information in question. The underlined font highlights relating information that has not been unlearned in the response, resembling to the ground truth.
@@
Table 11: Examples of responses after unlearning on TOFU (Retain95) against QAs targeted for unlearning. The bold font highlights the key information in question. The underlined font highlights deviate information that not consistent with the ground truth.

### v3

@@ B.5. Results on LLM unlearning
Figure 5. Comparison of forget quality and model utility between GA, NPO, SimNPO, and IMU.
B.6. Changes of influence weights over epochs
In Figure 6, we present the changes in the influence weights of $15$ samples randomly selected from $\mathcal{D}_{f}$ across $5$ unlearning epochs for both the class-wise and sample-wise unlearning tasks. In the class-wise unlearning task, since all samples belong to the same class, their influence values are negative according to Section 3.2. As shown in Figure 6, the weight of each sample stabilizes after approximately two epochs, and the gradual convergence of all weights toward similar contributions indicates that the model is effectively forgetting the high influence samples. In contrast, in the sample-wise unlearning scenario, the $15$ samples are randomly selected from different classes, and thus their overall influence on the forget set can be either positive or negative. As shown in Figure 6, the number of samples with negative influence values gradually increases over the unlearning process, suggesting that the model updates are progressively aligned with the direction that promotes forgetting, demonstrating the effectiveness of IMU.
Figure 6. Change of sample influences for $5$ epochs in CIFAR-10 class-wise and sample-wise unlearning tasks.
Table 14. Examples of responses after unlearning on TOFU (Forget05) against QAs targeted for unlearning. The bold font highlights the key information in question. The underlined font highlights relating information that has not been unlearned in the response, resembling to the ground truth.
@@ B.6. Changes of influence weights over epochs
Table 15. Examples of responses after unlearning on TOFU (Retain95) against QAs targeted for unlearning. The bold font highlights the key information in question. The underlined font highlights deviate information that not consistent with the ground truth.